1. 5.00 g NaOH and 5.00 g Ba(OH)2 were used to prepare 1L solution each. Which solution will be more basic? (MM NaOH=40, Ba(OH)2=171.3)



2. You decide to test the pH of your brand new swimming pool. The instruction manual advises to keep it between 7.2-7.6. Shockingly, you found out that the pH of your pool is 8.3! What kind of chemical should you add?




1
Expert's answer
2022-06-02T10:14:03-0400

1.

To determine which one is more basic, we first find the pOH of both solutions. whichever has a lower pOH solution a more basic.

n[NaOH] = 5 / 40 = 0,125 mol

CM (OH-)= n/V = 0,125/1 = 0,125 mol/l

pOH = -log [OH-] = -log (0,125) = 0,9



n[Ba(OH)2] = 5 / 171 = 0,029 mol

CM (OH) = 0,029 * 2 / 1 = 0,058 mol/l

pOH = -log [OH-] = -log (0,058) = 1,24


ANSWER: NaOH solution more basic


2.

pH = 7 — solution is neutral

pH < 7 — solution is acidic

pH > 7 — solution is basic

hence the basics increased in the example above. we need to add salts consisting of strong acids and weak bases to normalize it. For example ZnSO4 , AlCl3 ...